You’re given the following data concerning Freedonia, a legendary country:

1)     Consumption function: C=200+0.8Y

2)     Investment function: I=100

3)     AE=C+I

4)     AE=Y

a.      What is the marginal propensity to consume in Freedonia, and what is the marginal propensity to save?

b.     Graph equations (3) and (4) and solve for equilibrium income.

c.      Suppose aquation (2) is changed to (2’) I=110. What is the new equilibrium level of income? By how much does the $10 increase in planned investment change equilibrium income? What is the value of multiplier?

d.     Calculate the saving function for Freedonia. Plot this saving function on a graph with equation (2). Explain why the equilibrium income in this graph must be the same as in part b.



1
Expert's answer
2021-10-11T08:53:40-0400

a. The MPC is 0.8

and MPS is 0.2, as we know MPC+MPS=1

So MPS= 1-0.8=0.2

b. Y= 200+0.8Y+100=300 +0.8Y

Y-0.8Y=300

0.2Y=300

Y=300/0.2=1500

C=Y-I=1500-100=1400



c. If I=110

Y=C +I=200+0.8Y +110=310+0.8Y

Y-0.8Y=310

0.2Y=310

Y=310/02=1550

Thus, $10 increase in investment leads to $50 increase in income-consumptionMultiplier= Change in income/Change in Investment=50/10=5

d. Y=C+S

S=Y-200-0.8Y

=-200+0.2Y

The eqvilibrium is whese Saving=Investment